An elevator weighing 4000 Ib attains an upward velocity of 4 m/sec in 3 sec withuniform acceleration. Find the apparent weight of a 40 kg man standing inside the elevator during its ascent


1
Expert's answer
2021-06-23T13:02:44-0400

"N-mg=ma\to N=m(g+a)"


"v=at\to a=v/t=4/3\approx1.33\ (m/s^2)"


"N=m(g+a)=40\cdot(9.81+1.33)=445.73\ (N)" . Answer